A train of mass m moves with a velocity v on the equator from east to west. If w is the angular speed of the earth about its axis and R is its radius(of the earth) then the normal reaction acting on the train is-------------

Grade:12

2 Answers

vikas askiitian expert
13 years ago

relation bw gravitational field (g) & rotation of earth is given by

 g1 = g - W2RCOS2@

at equator , @ = 0 so

 g1 = g - W2R

normal reaction on train which is at equator is

N = mg1 = m(g-w2R)

Arjun
10 years ago
mg[1 - {(wR- 2v)w}/g - v^2/Rg]